# Fully Dynamic Matching and Ordered Ruzsa-Szemerédi Graphs

### v2

@@ 3 Preliminaries
A fully dynamic graph $G=(V,E)$ is a graph defined on a fixed vertex set $V$ that is subject to edge insertions and deletions. We assume that each edge update is issued by an adaptive adversary that can see our algorithm’s previous outputs and can accordingly decide on the next update. We say an algorithm has amortized update-time $U$ if the total time spent after $T$ updates is $U\cdot T$ for some sufficiently large $T=\operatorname{poly}(n)$.
@@
Proposition 3 ([31]).
Let $G=(V,E)$ be any (possibly non-bipartite) $n$-vertex graph. Suppose that for some parameters $\varepsilon\in(0,1)$ we have an algorithm $\mathcal{A}$ that provided any vertex subset $U\subseteq V$ with $\mu(G[U])\geq\varepsilon n$, finds a matching of size at least $\varepsilon\mu(G[U])$ in $G[U]$. Then there is an algorithm that finds a matching of size at least $\mu(G)-O(\varepsilon n)$ in $G$ by making $\widetilde{O}_{\varepsilon}(1)$ adaptive calls to $\mathcal{A}$ where preparing the subset $U$ for each call can be done in $\widetilde{O}_{\varepsilon}(n)$ time.
We also use the following sublinear-time algorithm of Behnezhad [6] for estimating the size of maximum matching. We note that even though we use this algorithm, we maintain the edges of the maximum matching explicitly.
@@
Finally, we use the following vertex sparsification idea which was adapted to the dynmaic setting in the work of Kiss [29].
Definition 5.
Let $V$ be a vertex set and denote $k\leq|V|/2$ be a parameter. A set $\mathcal{P}=\{P_{1},\ldots,P_{\ell}\}$ of $\ell$ partitionings of $V$ is a $(\ell,k,\varepsilon)$-vertex sparsifier if the following hold:
  1. 1.
    Each partitioning $P_{i}$ partitions $V$ into $O(k/\varepsilon^{2})$ subsets.
  2. 2.
    For each subset $S\subseteq V$ of size $2k$, there is at least one partitioning $P_{i}$ such that there are at least $(1-\varepsilon)|S|$ subsets in $P_{i}$ that have exactly one vertex of $S$.
Proposition 6 ([29]).
Let $\mathcal{P}$ include $\ell=\lceil 512\log n/\varepsilon^{2}\rceil$ random partitionings $P_{1},\ldots,P_{\ell}$ of $V$, where each $P_{i}$ partitions $V$ into at least $8k/\varepsilon$ subsets randomly. Then, with probability $1-1/\operatorname{poly}(n)$, $\mathcal{P}$ is a $(\ell,k,\varepsilon)$-vertex sparsifier of $V$.
The notion of graph contraction here refers to the process of simplifying the graph based on the partitionings provided by the vertex sparsifiers. Each set of partitions effectively represents a contracted version of the graph, where the vertices within the same subset can be considered as a single entity for the purpose of matching. This contraction not only reduces the overall size of the graph but also retains the essential structural properties that influence the size and composition of maximum matchings.
Definition 7.
Denote $M$ be a given matching with $2r$ vertices in graph $G$. Let us assume the $d_{v}$ for a vertex $v$ in the matching is the number of neighbors of $v$ in $V(M)$. we define average degree of $M$ as $\sum_{v\in V(M)}d_{v}/2r$.
4 Dynamic Approximate Matching
We present our dynamic matching algorithm in this section and prove Theorem 1.
Let us start with an overview of how our proposed algorithm is built. We randomly sample edges from graph $G$ to find a matching with $\varepsilon$-approximation, which is described in Section 4.1. We propose that if we need many samples to find this matching then the average degree of this matching is small in the remaining edges (what we mean by the remaining edges is that, when we output a matching, the adversary can remove its edges from the graph, while searching for a new matching we should discard the edges found before), resulting in the matching to be induced in the remaining graph. In this case, we charge the cost of the algorithm to ORS, with characteristics similar to the matching we found. Otherwise, we will succeed in finding a matching, relatively fast.
Note that, in this model, we can argue that if the matching size in the graph is $\mu$ then the matching size will not change after $\varepsilon^{2}n$ updates. We will find a matching once in $\varepsilon^{2}n$ edge updates to optimize amortized running time. This motivates us to run Section 4.1 to find a matching after $\varepsilon^{2}n$ updates. To keep track of the large matchings we already found we introduce Section 4.2. This algorithm simplifies our problem where the matching size does not change. If we maintain this algorithm for different matching sizes, knowing the approximate matching we will know which maintained algorithm to look at to find the $\varepsilon$-approximate matching.
To boost our approximation to $(1-\varepsilon)$ we use Proposition 3 to output the subsets of $U$ that we need to search for the matchings of size $\varepsilon\mu$. We use the subsets in Section 4.1 which is a $\varepsilon$-approximation algorithm.
Note that we provide a multiplicative approximation maximum matching algorithm for the dynamic model, which leads to the fact that we need to handle the cases where the matching size is not of order $n$. This adds some layers to the intuition we stated above.
  1. 1.
    If the matching size is $n^{\prime}$ we output a matching after $\varepsilon^{2}n^{\prime}$ updates. This means we need to handle matching with similar sizes together, we will have $\log n$ partition to keep track of the matchings outputted by Section 4.1 of almost the same size simultaneously.
  2. 2.
    To handle the graph with small matchings of $O(n^{\prime})$ we need to contract the graph to one with $n^{\prime}$ vertices. We build a contracted graph by creating $n^{\prime}$ buckets of $O(n/n^{\prime})$ vertices, the vertices are distributed randomly in buckets. Since we want to preserve the edges of the maximal matching we need to randomly maintain $O(\log n)$ such contractions for $\log n$ values of $n^{\prime}$.
  3. 3.
    The fact that the edges of the matching found in the earlier steps of Section 4.2, could be removed by the adversary leads us to a definition of the degree of matching that only takes into account the matchings that we will be finding in the next rounds (we do not care about the edges removed from the graph in the dynamic steps). This concept leads us to define the set of matchings named ORS, where every matching $M$ is induced in the set of matchings we observe after $M$.
4.1 Random Sampling Algorithm
Here we introduce a sampling algorithm that given a graph $G$, with a matching of size $\mu$ can output a matching of size $\varepsilon\mu$. We bound the algorithm running time with the average degree of the outputted matching. The way the algorithm works is that for each vertex we sample $b$ vertices and check if the edge is in the graph using the adjacency matrix of the graph. We add the edges to the matching if none of the endpoints have been used previously. Note that by starting with a small $b$ for the budget, we ensure to find the edges from dense parts of the graph which is easier to find. As we go further in the algorithm we make the budget larger to find edges in more sparse parts of the graph. Now, we state the random sampling algorithm and prove its efficiency.
@@ 4.1 Random Sampling Algorithm
Given a graph $G=(V,E)$, vertex subset $U\subseteq V$, and parameters $\varepsilon\in(0,1)$, such that maximal matching of $G[U]$ is $\mu\geq\varepsilon n$, there is an algorithm that finds a matching $M\subseteq G[U]$ of size at least $\varepsilon\mu$. The algorithm runs in time $O(n^{2+\varepsilon}/d)$ where $d=|E\cap V^{2}(M)|/O(n)$.
@@
Let us consider Section 4.1, first we prove that this algorithm will provide a matching $M_{c}$ of size a least $\varepsilon\mu/2$ by sampling vertices with budget $b\in[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$. Let us assume there are no such intervals. This means that for any integer $c\in[0,1/(2\varepsilon)]$, running the algorithm for $b$ in the interval $[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$ outputs a matching with less than $\varepsilon\mu$ edges. Note that for each edge $e$ in $M_{1}\cup\dots\cup M_{1/(2\varepsilon)}$, at most 2 edges in the maximum matching of $G[U]$ have common vertices with $e$. Summing over all the $1/(2\varepsilon)$ intervals we have:
$22\sum_{c=1}^{1/(2\varepsilon)}{|M_{i}|}<2\cdot\frac{1}{2\varepsilon}\cdot\varepsilon\mu/2=\mu/$ 2 .
The fact that we could not find a matching of size $\mu/2$ is a contradiction with the lemma statement ensuring that the size of maximum matching is at least $\mu/2$. Note that by exhausting the algorithm with budget of $n$, we will go through all the edges and output a maximal matching. Note that any maximal matching in this graph should be of size at least $\mu/2$.Therefore, we have at least the size of one of the matchings in$\{M_{1},...,M_{1/(2\varepsilon)}\}$ is $\varepsilon\mu/2$.
Now, we prove the statement of the algorithm. We assume that this algorithm terminates at $b$ such that $2^{\varepsilon(c-1)\cdot\log n}\leq b<2^{\varepsilon c\cdot\log n}$, for a constant $\delta$, we spend $n[b+b/2+..+1]<\delta nb\leq\delta n\cdot 2^{\varepsilon c\log n}<\delta n^{1+\varepsilon c}$ time to find the matching $M_{c}$ such that $|M_{c}|\geq\varepsilon\mu/2$. Note that if it takes at most $n^{2+\varepsilon}/d$ time to find the matching of size $\varepsilon\mu/2$, then we have $n^{2+\varepsilon}/d<\delta n^{1+\varepsilon c}$ therefore, $d>\frac{n^{1+\varepsilon-\varepsilon c}}{\delta}\geq\frac{n^{1+\varepsilon}}{\delta}$.
For budget $b$, we sample vertices for all the vertices, we argue that for the vertices that are not in $M$, and after sampling $b$ vertices we could not find their neighbor then the degree of vertices out of the matching is less than $O(n/b)$. Assuming that the degree of vertex $v$ is $d_{v}=c\cdot\frac{n}{b}$, with probability $(1-\frac{d_{v}}{n})^{b}=(1-c\cdot\frac{1}{b})^{b}\approx(1/e)^{c}$ we sampled a neighbor. Therefore in this case with a constant probability we already matched $v$. This implies that if we find the first edge of this matching at a budget of $2b$, we get that the degree of any vertex out of the matching we found until this point is less than $c\cdot\frac{n}{b}$ for any constant $c$. Picking $c$ small enough, the total sum of the degree of the matching is at most $\varepsilon\mu c\cdot\frac{n}{b}<\varepsilon\mu n^{(1-\varepsilon(c-1))}=O(nd)$. ∎
4.2 The Overall Algorithm via Certificates and Caching Them
In the next step, we aim to use Section 4.1 efficiently to output an approximate matching in the dynamic model. First, we prove that given the matching size is equal to $O(n^{\prime})$, after $\varepsilon^{2}n^{\prime}$ updates the matching size does not change substantially. This means we can output a matching after each $\varepsilon n^{\prime}$ update and get an $\varepsilon$-approximate matching.
Observation 9.
Given a graph $G=(V,E)$ with $n$ vertices, if the maximum matching size is $O(\mu)$ and we provide an $\varepsilon$-approximate matching $M^{\prime}$, after $\varepsilon^{2}\mu$ updates in the dynamic model, $M^{\prime}$ or one of its subsets remains an $\varepsilon$-approximate matching for the updated graph.
@@ 4.2 The Overall Algorithm via Certificates and Caching Them
We assume that in a graph $G$ the maximal matching size is $c\mu$ where $c$ is a constant, after $\varepsilon^{2}\mu$ updates in the graph the size of the maximal matching is between $(c-\varepsilon^{2})\mu,(c+\varepsilon^{2})\mu$. This implies that if we have an algorithm that outputs an $\varepsilon c\mu$ matching, after $\varepsilon^{2}\mu$ updates, the $\varepsilon$ approximate matching is preserved. This is because we lose at most $\varepsilon^{2}$ fraction of the outputted matching. ∎
Note that to use Section 4.1 we need to know about the existence of a matching $\mu$, the algorithm described in Proposition 4, given a graph with $n$ vertices, outputs if there is matching of size $O(n)$. Using this algorithm alone will not help us with the cases where the matching size is not of order $n$. This motivates us to find an algorithm that reduces the number of vertices on the graph while maintaining the maximum matching size. Proposition 6 will maintain $O(\log n)$ contraction on the graph for an integer $k\in[1,\log n]$, such that for any matching $M$ with high probability, there is a contraction preserving a $1-\varepsilon$ approximation of $M$. After running the reduction on the graph, we find the contraction that preserves the maximum matching. Let us assume that $G^{\prime}$ is the contracted graph, we run Section 4.1 on $G^{\prime}$ to output the $\varepsilon$ approximation matching.
Note that throughout the changes in the dynamic model, both the matching size and the contracted graph that preserves the maximum matching, change. An edge insertion in $G$ is a weight addition to an edge in graph $G^{\prime}$, and edge deletion is a weight subtraction in $G^{\prime}$. In the following algorithms, when we find a matching in $G^{\prime}$, we do not care about the weight of edges, an edge in $G^{\prime}$ exists if its weight is more than 0, otherwise, This edge does not exist in $G^{\prime}$.
First, we solve a simplified version of the problem where the matching size does not change throughout the updates, and a single contracted graph $G^{\prime}$ preserves matching edges in the dynamic model. We prove that there exists an efficient algorithm maintaining a $(1-\varepsilon)$-approximation matching in the dynamic model.
We introduce Section 4.2. We assume that the updates in this algorithm are in the case that the matching size in $O(n^{\prime})$, and $G^{\prime}$ preserves the maximum matching of the graph. This algorithm works in phases, each phase represents the number of updates we want to use the matchings we found in the last updates.

### v3

@@ 3 Preliminaries
A fully dynamic graph $G=(V,E)$ is a graph defined on a fixed vertex set $V$ that is subject to edge insertions and deletions. We assume that each edge update is issued by an adaptive adversary that can see our algorithm’s previous outputs and can accordingly decide on the next update. We use $\mu(G)$ to denote the maximum matching size in G. We say an algorithm has amortized update-time $U$ if the total time spent after $T$ updates is $U\cdot T$ for some sufficiently large $T=\operatorname{poly}(n)$.
@@
Proposition 3 (Approximation Boosting Framework [37, 24]).
Let $G=(V,E)$ be any (possibly non-bipartite) $n$-vertex graph. Suppose that for any parameter $\delta\in(0,1)$ we have an algorithm $\mathcal{A}(G,U,\delta)$ that provided any vertex subset $U\subseteq V$ with $\mu(G[U])\geq\delta\cdot n$, finds a matching of size at least $\operatorname{poly}(\delta)\cdot\mu(G[U])$ in $G[U]$. Then for any $\varepsilon\in(0,1)$, there is an algorithm that finds a matching of size at least $\mu(G)-\varepsilon n$ in $G$ by making $t=O_{\varepsilon}(1)$ adaptive calls
$\mathcal{A}(G,U_{1},\delta_{1}),\ldots,\mathcal{A}(G,U_{t},\delta_{t})$
to algorithm $\mathcal{A}$. The value of $\delta_{i}$ in each of these calls is just a function of $\varepsilon$, $\delta_{i}\leq\varepsilon/2$, and preparing the vertex subsets $U_{1},\ldots,U_{t}$ can be done in $\widetilde{O}_{\varepsilon}(n)$ total time.
We also use the following sublinear-time algorithm of Behnezhad [10] for estimating the size of maximum matching. We note that even though we use this algorithm in a crucial, our final dynamic algorithm does not maintain just the size, but rather the edges of the matching, explicitly.
@@
Finally, we use the following proposition which turns an additive approximation into a multiplicative approximation. The proof is based on a vertex sparsification idea for matchings [5, 14] which was adapted to the dynamic setting in the work of Kiss [34].
Proposition 5 ([34, 5]).
Suppose there is an adaptive algorithm $\mathcal{A}$, that for any parameter $\varepsilon>0$ and a fully dynamic $n$-vertex graph $G=(V,E)$, maintains a matching of size $\mu(G)-\varepsilon n$ in $Q(n,\varepsilon)$ amortized time. Then there is an algorithm that maintains a multiplicative $(1-\varepsilon)$-approximation maximum matching of $G$ in $\operatorname{poly}(\log(n),\varepsilon)\cdot Q(n,\varepsilon^{2})$ amortized time per update.
4 Our Dynamic Algorithm for Approximate Maximum Matching
In this section, we present our algorithm and prove Theorem 1. We start in Section 4.1 with a static potentially sublinear-time algorithm that is one of the main building blocks of our final algorithm. We then formalize our dynamic algorithm in Section 4.2. In Section 4.3 we prove correctness of our dynamic algorithm and analyze its running time in Section 4.4.
4.1 A Static Potentially-Sublinear Time Algorithm via Random Sampling
In this section, we present a potentially sublinear time algorithm that given an $n$ vertex graph $G=(V,E)$, finds a matching $M$ that is by an additive factor of at most $\varepsilon n$ smaller than the maximum matching of $G$. The algorithm, in addition, returns a certificate $M_{C}$, which is another matching of $G$ that we use to bound the running time of the algorithm. Particularly, denoting by $d$ the average degree in the subgraph induced on the vertices of the certificate (i.e., graph $G[V(M_{C})]$), the running time of the algorithm overall is $\widetilde{O}(n^{2+2\varepsilon}/d)$. If the certificate $M_{C}$ is close to an induced matching (i.e., $d$ is small), then the algorithm is not better than the trivial algorithm which reads all the edges in quadratic in $n$ time. On the other hand, if $G[V(M_{C})]$ is dense, the algorithm runs in sublinear time. This algorithm will be a crucial component of our final dynamic algorithm.
The following is the main lemma of this section.
@@ 4.1 A Static Potentially-Sublinear Time Algorithm via Random Sampling
Let $G=(V,E=E_{\text{dense}}\cup E_{\text{sparse}})$ be a given $n$-vertex graph where we have adjacency matrix access to $E_{\text{dense}}$, adjacency list access to $E_{\text{sparse}}$, and $E_{\text{dense}}$ and $E_{\text{sparse}}$ may share edges. For any parameter $\varepsilon\in(0,1)$, there is an algorithm $\textsc{MatchAndCertify}(E_{\text{dense}},E_{\text{sparse}},\varepsilon)$ that returns a $(1,\varepsilon n)$-approximate maximum matching $M$ of $G$ and a certificate $M_{C}$. Let $T$ be the running time of the algorithm. Then exactly one of the following two conditions holds:
  1. $(C1)$
    $M_{C}=\emptyset$ and $T=\widetilde{O}_{\varepsilon}(n+|E_{\text{sparse}}|)$.
  2. $(C2)$
    $M_{C}$ is a matching in $E_{\text{dense}}$ of size $\Omega_{\varepsilon}(n)$ where $M_{C}\cap E_{\text{sparse}}=\emptyset$. Additionally, $T=\widetilde{O}_{\varepsilon}(|E_{\text{sparse}}|+n^{2+\varepsilon}/d)$ where $d:=|E_{\text{dense}}\cap V^{2}(M_{C})|/|M_{C}|$ is the average degree of $E_{\text{dense}}[V(M_{C})]$.
Towards proving Lemma 6, we first prove the following Lemma 7 which solves a slightly simpler subproblem that finds a large matching in a given induced subgraph of $G$. We will later prove Lemma 6 by combining Lemma 7 with Proposition 3.
Lemma 7.
Let $G=(V,E)$ be a given $n$-vertex graph to which we have adjacency matrix access. Let $U\subseteq V$ be a given vertex subset, and let $\delta\in(0,1)$ be a given parameter such that $\mu(G[U])\geq\delta n$. There is an algorithm that finds a matching $M\subseteq G[U]$ of size at least $\delta^{2}n$. The algorithm runs in $\widetilde{O}(n^{2+2\delta}/d)$ time where $d=|E\cap V^{2}(M)|/|M|$.
@@ 2. $(C2)$
Note that if the algorithm returns a matching $M_{i}$, then it is guaranteed that the matching has our desired size of at least $\delta^{2}n$. So it suffices to prove that at some point $M_{i}$ is as large as we want. To see this, note that once the budget $b$ reaches $n$, we will go through all the edges, indicating that $M$, the union of matching $M_{1},M_{2},\dots,M_{1/(2\delta)}$, is a maximal matching of $G[U]$. Since the size of a maximal matching is at least half the size of a maximum matching, we have $|M|\geq\mu(G[U])/2\geq\delta n/2$. This means that at least one of the matchings in $\{M_{1},...,M_{1/(2\delta)}\}$ is of size at least $\frac{\delta^{2}n/2}{1/(2\delta)}=\delta^{2}n$.
Next, we focus on relating the running time of the algorithm to the parameter $d$, i.e., the average degree of the output matching of Section 4.1. Let us first introduce some notation. Let $U^{\prime}_{i}$ be the value of the set $U^{\prime}$ at the end of iteration $i$. Additionally, for every vertex $v$, let $d^{i}_{v}$ be the degree of vertex $v$ in graph $G[U^{\prime}_{i}]$. We start by proving the following useful claim.
Claim 8.
It holds with probability $1-1/n$ for every $v\in U^{\prime}_{i}$ and every $i$ that $d^{i}_{v}=O(n^{1-2\delta(i-1)}\log n)$.
@@
Fix $v$ and $i$. We prove that either $v\not\in U^{\prime}_{i}$ or $d^{i}_{v}=O(n^{1-2\delta(i-1)}\log n)$ with probability at least $1-1/n^{3}$. A union bound over the choices of $v$ and $i$ completes the proof.
Consider iteration $i$ of the algorithm and suppose $v\in U^{\prime}_{i}$. This means that in this iteration we attempt to match vertex $v$ by sampling $b=n^{2\delta(i-1)}$ vertices in $U^{\prime}\supseteq U^{\prime}_{i}$ and checking if any of them is a neighbor of $v$. Let $d^{\prime\prime}$ be the number of neighbors of $v$ in $U^{\prime}$ when processing $v$ in iteration $i$. We have
$\Pr[v\text{ not matched by }M_{i}]\leq\left(1-\frac{d^{\prime\prime}}{|U^{\prime}|}\right)^{b}\leq\left(1-\frac{d^{\prime\prime}}{n}\right)^{b}\leq\exp\left(-\frac{d^{\prime\prime}b}{n}\right)=\exp\left(-\frac{d^{\prime\prime}}{n^{1-2\delta(i-1)}}\right).$
Now consider two cases. If $d^{\prime\prime}\geq 3n^{1-2\delta(i-1)}\log n$, then $v$ is matched in $M_{i}$ with probability at least $1-1/n^{3}$ and thus will be removed from $U^{\prime}$. Otherwise, since $d^{i}_{v}\leq d^{\prime\prime}$, we have $d^{i}_{v}=O(n^{1-2\delta(i-1)}\log n)$. Thus, in either case, the statement holds. ∎
Let $M_{i}$ be the matching that Section 4.1 returns. Note that in iteration $j$ we have $b=n^{2(j-1)\delta}$ and the algorithm takes $O(nb)=O(n^{1+2(j-1)\delta})$ time. So the running time of the algorithm until termination is $O(n)+O(n^{1+2\delta})+\ldots+O(n^{1+2(i-1)\delta})=O(n^{1+2(i-1)\delta})$. Since every vertex $v$ matched by $M_{i}$ must belong to $U^{\prime}_{i-1}$, applying Claim 8 gives $d^{i-1}_{v}=O(n^{1-2\delta(i-2)}\log n)$. This implies that the total number of edges in $G[V(M_{i})]$ can be upper bounded by $|M|\cdot O(n^{1-2\delta(i-2)}\log n)$. Letting $d=\Theta(n^{1-2\delta(i-2)}\log n)$, the running time can be re-written as
$O(n^{1+2(i-1)\delta})=O(n^{1+2(i-1)\delta}d/d)=O(n^{1+2(i-1)\delta+1-2\delta(i-2)}\log n/d)=O(n^{2+2\delta}\log n/d),$
which is the claimed bound. ∎
Next, we apply Proposition 3 on Lemma 7 to complete the proof of Lemma 6.
Proof of Lemma 6.
For any graph $H$, let GreedyMatching($H$) be a greedy algorithm that goes over all the edges of $H$ one by one and adds an edge $(u,v)$ to the matching if $u$ and $v$ are not matched already. Note that the output of this algorithm is a maximal matching in $H$.
The algorithm MatchAndCertify proceeds as follows. Our goal is to run the algorithm of Proposition 3 to find a matching of size $\mu(G)-\varepsilon n$ in $G$. This algorithm makes $t=O_{\varepsilon}(1)$ adaptive calls
$\mathcal{A}(G,U_{1},\delta_{1}),\dots,\mathcal{A}(G,U_{t},\delta_{t})$
to an algorithm $\mathcal{A}$ that outputs a matching of size $\operatorname{poly}(\delta_{i})\cdot\mu(G[U_{i}])$ in $G[U_{i}]$ provided that $\mu(G[U_{i}])\geq\delta_{i}n$. Here $U_{i}$ is a subset of $V$ and parameter $\delta_{i}\leq\varepsilon$ is a function of $\varepsilon$. In order to use Proposition 3, we have to specify how the algorithm $\mathcal{A}$ is implemented.
We implement $\mathcal{A}(G,U_{i},\delta_{i})$ as follows. Let $G_{i}:=G[U_{i}]\cap E_{\text{sparse}}$. As the first step, we call $\textsc{GreedyMatching}(G_{i})$ to obtain matching $M_{i}$ in $O(|E_{\text{sparse}}|)$ time. If $|M_{i}|\geq\delta^{2}_{i}\cdot n/8$, then we have a large enough matching in $G[U_{i}]$, and $\mathcal{A}$ terminates by outputting $M_{i}$ and we set $d_{i}=\infty$. Otherwise, we run $\textsc{RandomSampling}(G[U_{i}]\cap E_{\text{dense}},\delta_{i})$ from Lemma 7. Note that Lemma 7 requires adjacency matrix access to its input graph which we can provide, given the guarantee of the lemma that we have adjacency matrix access to $E_{\text{dense}}$ and that the set $U_{i}$ is given explicitly. Since GreedyMatching returns a 1/2-approximate maximum matching, we get that
$\mu(G[U_{i}]\cap E_{\text{dense}})\geq\mu(G[U_{i}])-\mu(G[U_{i}]\cap E_{\text{sparse}})\geq\delta_{i}n-2|M_{i}|=3\delta_{i}n/4.$
Therefore, algorithm RandomSampling outputs a matching $M_{C_{i}}$ of size at least $169\delta^{2}_{i}n/$ 16 in $G[U_{i}]\cap E_{\text{dense}}$. By Lemma 7, this runs in $\widetilde{O}(n^{2+2\delta_{i}}/d_{i})=\widetilde{O}(n^{2+\varepsilon}/d_{i})$ time where $d_{i}=|E\cap V^{2}(M_{C_{i}})|/|M_{C_{i}}|$.
From our discussion above, in the $i$-th call to $\mathcal{A}$, this algorithm outputs a matching of size at least $\min\{\delta_{i}^{2}n/8,9\delta_{i}^{2}n/16\}=\delta_{i}^{2}n/8$. Therefore, the algorithm of Proposition 3 outputs a matching $M$ of size $\mu(G)-\varepsilon n$.
It remains to analyze the running time by specifying the outputs $d$ and $M_{C}$ and showing that at least one of the conditions $(C1)$ or $(C2)$ hold.
If we never call the RandomSampling algorithm, we simply set $d=0$ and $M_{C}=\emptyset$ as all calls to $\mathcal{A}$. The total running time spent by Proposition 3 preparing the vertex subsets $U_{i}$ for each $i\in[t]$, and the running time of the calls to the GreedyMatching is
$\widetilde{O}_{\varepsilon}(n)+\sum_{i\in[t]}O(|E(G_{i})|)=\widetilde{O}_{\varepsilon}(n)+t|E_{\text{sparse}}|=\widetilde{O}_{\varepsilon}(n+|E_{\text{sparse}}|).$
Thus, condition $(C1)$ holds.
Otherwise, we let $j=\operatorname*{arg\,min}_{i\in[t]}d_{i}$, set $d:=d_{j}$ and return the certificate $M_{C}=M_{C_{j}}\setminus E_{\text{sparse}}$. We now show that condition $(C2)$ holds.
Let us now bound the running time. This is the cost of preparing the vertex subsets of Proposition 3, the running time of the calls to the GreedyMatching, and the running time of the calls to RandomSampling. The total running time is therefore
$\widetilde{O}_{\varepsilon}(n)+t|E_{\text{sparse}}|+\widetilde{O}\left(\sum_{i=1}^{t}\frac{n^{2+\varepsilon}}{d_{i}}\right)=\widetilde{O}_{\varepsilon}\left(|E_{\text{sparse}}|+\frac{n^{2+\varepsilon}}{d}\right).$
The bound above is the one required by $(C2)$, except that we defined $d=d_{j}=|E\cap V^{2}(M_{C_{j}})|/|M_{C_{j}}|$ whereas the certificate $M_{C}$ is a subset of $M_{C_{j}}$. To fix this, in what follows we prove that $d$ is at least a third of the value of the right parameter $d^{\prime\prime}:=|E\cap V^{2}(M_{C})|/|M_{C}|$, which completes the proof.
The fact that we run RandomSampling on $G[U_{j}]$ implies that $\textsc{GreedyMatching}(G_{j})$ did not return a large enough matching. This implies that $\mu(G_{j})\leq\delta^{2}_{j}n/4$. Therefore, once we remove the edges of $E_{\text{sparse}}$ from $M_{C_{j}}$, we reduce the size of this matching by at most $\delta^{2}_{j}n/4$. Combined with $|M_{C_{j}}|\geq 9\delta_{i}^{2}n/16$, this implies that $|M_{C}|\geq|M_{C_{j}}|-\delta_{j}^{2}/4\geq 7|M_{C_{j}}|/16$. Hence
$d^{\prime\prime}=\frac{|E\cap V(M_{C})|}{|M_{C}|}\leq\frac{|E\cap V(M_{C_{j}})|}{7|M_{C_{j}}|/16}<3d,$
completing the proof as discussed above. ∎
4.2 Our Dynamic Algorithm via Lemma 6
In this section, we formalize our dynamic algorithm for Theorem 1 based on the static MatchAndCertify algorithm of Lemma 6. Our algorithm is formalized as Section 4.2.
We prove in Section 4.3 that our algorithm maintains an additive $(1,\varepsilon n)$-approximation of the graph at anytime.
Lemma 9 (Correctness of Section 4.2).
At any point, the output matching $M_{G}$ of Section 4.2 is a $(1,\varepsilon n)$ approximate maximum matching of graph $G$.
Later in Section 4.4, we analyze the running time of the algorithm and prove the following bound on it.
Lemma 10 (Runtime of Section 4.2).
Section 4.2 runs in time
$\widetilde{O}_{\varepsilon}\left(\frac{t}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}{t}\right),$
where $t$ is the threshold of the algorithm.
The combination of the two lemmas above with Proposition 5 implies Theorem 1.
Proof of Theorem 1.
The running time of Section 4.2 is analyzed in Lemma 10, which gives a boundo f
$\widetilde{O}_{\varepsilon}\left(\frac{t}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}{t}\right),$
where $t$ is a threshold parameter. To achieve the amortized time, we balance the two terms in the running time expression by setting
$t=\sqrt{n^{3+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}.$
By substituting this value of $t$ we have
$\widetilde{O}_{\varepsilon}\left(\frac{\sqrt{n^{3+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}{\sqrt{n^{3+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}}\right)=\widetilde{O}_{\varepsilon}\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}\right).$
By Lemma 9, Section 4.2 maintains a matching of size $\mu(G)-\varepsilon n$ at any point. By applying Proposition 5, we can boost this additive approximation to a multiplicative $(1-\varepsilon)$-approximation while keeping the update time $\widetilde{O}_{\varepsilon}\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}\right).\qed$